Known H2CO3: = 4.30×10-7, = 5.61×10-11, the pH of 0.1mol·L-1 sodium bicarbonate solution is ?
1
Expert's answer
2021-01-13T13:54:59-0500

pH is the negative logarithm to base 10 of the hydrogen ion concentration.

Similarly, pOH is the negative logarithm to base 10 of the hydroxonium ion concentration.

Also, pKw of water is 14 at 25°C, therefore pH+pOH=14.

For buffer solutions:

pH = pKa + "Log \frac{[A^-]}{[HA]}" Where Ka is the activity of the acid, [A] is the salt concentration and [HA] the acid concentration. This equation is known as the Henderson equation.

Therefore, for the case of bicarbonate

pOH=-log(0.1)=1

But pH+pOH=14

Therefore, pH=14-pOH

pH=14-1=13